To determine the number of bass in a lake, a conservationist catches 54 bass, tags them, and returns them to the lake. Later, 32 are observed and 8 are tagged. Estimate how many bass are in the lake.
Answer by stanbon(75887)

To determine the number of bass in a lake, a conservationist catches 54 bass, tags them, and returns them to the lake. Later, 32 are observed and 8 are tagged. Estimate how many bass are in the lake.
-----------------------------------------------
[pop #/caught & tagged] = [caught including tagged]/[# tagged that were caught]
----------------------------
pop count/54 = 32/8
---
population count = 54[32/8) = 54*4 = 216
----------------------------------------------
Cheers,
Stan H.